OpenStudy (anonymous):

If f(x)=-1 what is x?

OpenStudy (agent0smith):

x could be any number, without more info.

OpenStudy (anonymous):

sorry it didnt let me finish lol ... f(x)=2x^2-x-1 ....if f(x) = -1 what is x?

OpenStudy (agent0smith):

That means 2x^2-x-1 = -1 first add 1 to both sides... 2x^2-x-1 +1 = -1 + 1 which simplifies to what?